# The Evolution of Dataset Distillation: Toward Scalable and Generalizable Solutions

### v2

@@ 3.4 Other Applications
Object Detection and Super-Resolution For object detection, Qi et al. [171] introduced DCOD, the first dataset condensation framework for this task, delivering competitive results on benchmarks like Pascal VOC [172] and MS COCO [173]. Dietz et al. [174] demonstrate the first comprehensive exploration of dataset distillation for image super-resolution (SR) [175], successfully reducing the original dataset size by 91.12% while maintaining comparable model performance through both pixel-space and latent-space distillation techniques, thereby offering a promising memory- and computation-efficient approach.
@@ 5 Challenges and Future Directions
Despite remarkable advancements in dataset distillation, critical challenges remain, particularly as the field shifts toward larger datasets and more complex applications.
@@ 5.1 Challenges
Scalability to Large-Scale Datasets While recent methods have demonstrated significant progress on ImageNet-1K, with approaches like D4M [61] achieving 68.1% accuracy at IPC=200 and CUDD [102] reaching 65.0%, scaling to larger datasets like ImageNet-21K remains challenging. On ImageNet-21K, even state-of-the-art methods like CUDD achieve only 34.9% accuracy at IPC=20, highlighting the substantial performance gap that exists when scaling to ultra-large datasets. These challenges stem from increased class counts, intra-class variability, and computational demands of large-scale datasets.
Balancing Compression and Performance The relationship between IPC and performance exhibits distinct patterns across different scales. While significant gains are observed between IPC=1 and IPC=10 (e.g., DATM [100] improves from 46.9% to 66.8% on CIFAR-10), performance improvements diminish at higher IPC values. For instance, on CIFAR-100, PAD [98] shows only marginal gains from IPC=50 (55.9%) to IPC=100 (58.5%). This non-linear relationship presents a critical trade-off between dataset compression and model performance, particularly crucial for practical applications.
Limited Cross-Architecture Generalization Current methods, particularly those based on SRe2L frameworks, often show architecture-specific performance patterns. While some methods demonstrate strong performance, their effectiveness across different architectures, especially emerging ones like Vision Transformers, remains less explored. The challenge of creating synthetic datasets that generalize well across diverse architectures persists.
Underexplored Domains and Modalities The majority of dataset distillation research has focused on vision tasks, with limited exploration in audio and multimodal domains. Emerging methods for temporal and multimodal data distillation demonstrate potential but require further development to match the maturity of image-based techniques.
Fair Evaluation of Dataset Distillation Methods Recent DD methods are shifting from hard labels to soft labels to enhance model accuracy [54]. However, soft labels inherently transfer additional knowledge from a teacher model, making it difficult to isolate the true informativeness of distilled data [179, 55]. Additionally, the use of diverse data augmentation techniques during evaluation further skews results, making it unclear whether performance gains stem from the quality of the distilled data or external enhancements [179]. Furthermore, different dataset distillation methods employ varied evaluation strategies, such as distinct loss functions, leading to inconsistencies in performance comparison.
@@ 5.2 Future Directions
Advanced Scaling Techniques Future research should focus on designing algorithms capable of scaling to ultra-large datasets without sacrificing performance. Techniques like hybrid approaches that combine distillation with coreset selection or data pruning could be explored to tackle computational bottlenecks.
Adaptive IPC Optimization Developing adaptive IPC optimization strategies that dynamically adjust IPC settings based on dataset complexity and downstream tasks could provide a practical solution to the compression-performance trade-off. Meta-learning and reinforcement learning frameworks offer promising avenues for automated IPC tuning.
Theoretical Understanding While dataset distillation methods have demonstrated impressive empirical success, their theoretical foundations remain relatively unexplored. A rigorous theoretical framework is needed to understand the fundamental limits of distillation, including convergence guarantees, optimal compression ratios, and generalization bounds. Such theoretical insights could provide principled guidance for designing more effective distillation algorithms and help understand the relationship between distillation size, model architecture, and performance guarantees.
Architecture-Agnostic Distillation Developing techniques that effectively transfer knowledge across different model architectures remains a critical challenge. Techniques that disentangle architecture-invariant features from architecture-specific biases could be further explored to enable broad applicability across architectures.
Domain Expansion Dataset distillation’s applicability should be extended to emerging domains, such as medical imaging, scientific analysis, and 3D point cloud data. Tailored frameworks that account for domain-specific characteristics, such as temporal dynamics in video or frequency-space representations in audio, are critical for broadening the scope of distillation techniques.
Enhancing Fair Evaluation and Comparisons In this survey, we present two comparative tables summarizing key aspects of various dataset distillation methods, including their use of soft labels, architectures, and other relevant factors, providing researchers with a reference to determine the most suitable approach for their needs. Additionally, Li et al. introduced DD-Ranking [179], a benchmark designed to decouple the effects of knowledge distillation and data augmentation, ensuring a fair evaluation framework to assess the true informativeness of distilled data. By combining structured comparisons with insights from existing benchmarks such as DD-Ranking [179], we aim to help researchers navigate the landscape of dataset distillation methods while promoting more standardized evaluation practices.
@@ Acknowledgment
The authors would like to express their gratitude to the contributors of the Awesome-Dataset-Distillation repository [^1] https://github.com/Guang000/Awesome-Dataset-Distillation, which has served as a valuable resource in compiling recent advancements.

### v3

@@ 3.2 Multi-modal Dataset Distillation
Most recently, Zhang et al. [184] provide a systematic diagnosis of why prior multimodal dataset distillation methods struggle. They identify a fundamental conflict between over-compression (inherent to dataset distillation) and contrastive supervision (essential for cross-modal alignment), which leads to modality collapse—a phenomenon where intra-modal representations become overly concentrated while inter-modal divergence increases. To address this, they propose RepBlend, a lightweight yet effective framework that mitigates modality collapse through representation blending to enhance intra-modal diversity and symmetric trajectory matching to balance supervision across modalities. Beyond performance gains, RepBlend also highlights future research directions such as fine-grained cross-modal alignment.
@@ 3.4 Other Applications
Object Detection and Image Restoration For object detection, Qi et al. [191] introduced DCOD, the first dataset condensation framework specifically tailored to this task, demonstrating competitive performance on standard benchmarks such as Pascal VOC [192] and MS COCO [193]. Extending dataset distillation beyond detection, recent approaches have also addressed challenging image restoration tasks. For instance, Dietz et al. [194] proposed a distillation framework focusing on adaptive super-resolution tasks, while Peng et al. [195] developed an instance-aware distillation method aimed explicitly at improving super-resolution performance. In parallel, Zheng et al. [196] introduced TripleD, a distribution-aware dataset distillation framework, explicitly designed to optimize image restoration tasks by better preserving underlying data distributions.
@@ 5 Challenges and Future Directions
Despite significant advancements in dataset distillation, several fundamental challenges persist, particularly as the field scales to larger datasets, diverse architectures, and multimodal applications. Addressing these challenges requires a combination of theoretical insights, scalable algorithms, and robust evaluation frameworks.
@@ 5.1 Challenges
Scalability to Large-Scale Datasets While recent dataset distillation methods have demonstrated notable progress on ImageNet-1K, further scaling to ultra-large datasets such as ImageNet-21K remains a major challenge. Even state-of-the-art methods like CUDD [115] achieve only 34.9% accuracy at IPC=20, revealing a substantial performance drop when applied to datasets with extensive class diversity and intra-class variability. Computational constraints, increased dataset redundancy, and high memory costs further complicate large-scale distillation.
Balancing Compression and Performance The trade-off between dataset compression and performance varies significantly across dataset sizes and IPC levels. While significant performance improvements occur at lower IPC values (e.g., DATM [113] improves from 46.9% to 66.8% on CIFAR-10 from IPC=1 to IPC=10), these gains diminish at higher IPCs. On CIFAR-100, PAD [111] shows only minor improvements from IPC=50 (55.9%) to IPC=100 (58.5%), suggesting a saturation point in dataset efficiency.
Cross-Architecture Generalization Current methods exhibit strong performance on pre-selected architectures but often fail to generalize across different model families. Vision Transformers and emerging architectures demonstrate different feature learning mechanisms from traditional CNNs, making it unclear whether existing distillation techniques transfer effectively.
Underexplored Domains and Modalities Most existing distillation methods are designed for vision tasks, particularly image classification, while other core vision domains, such as object detection and semantic segmentation, remain largely underexplored [200]. Beyond vision, distillation in other modalities such as audio [201], 3D point clouds [202], and multimodal learning [203] is still in its infancy. Challenges such as temporal consistency in video, frequency-space representations in audio, and cross-modal alignment in multimodal tasks are yet to be adequately addressed.
Standardizing Fair Evaluation Practices Recent methods increasingly rely on soft labels and teacher-student distillation strategies, making it difficult to separate the impact of distilled data from external knowledge transfer. Additionally, the use of varied evaluation protocols, loss functions, and augmentation strategies leads to inconsistencies in benchmarking, making cross-method comparisons less reliable.
@@ 5.2 Future Directions
Ultra-Scalable Dataset Distillation Achieving effective dataset distillation on billion-scale datasets like JFT-300M [204] and LAION-5B [205] requires new strategies. Techniques such as dataset pruning integrated with distillation, and self-supervised data compression hold promise for tackling scalability challenges while maintaining high downstream task performance.
Dynamic and Task-Aware Compression Instead of relying on fixed IPC values, future methods should explore dynamic IPC optimization tailored to dataset complexity and target tasks. Meta-learning and reinforcement learning frameworks could be leveraged to automatically tune IPC levels based on data redundancy and model requirements.
Towards Architecture-Agnostic Dataset Distillation Developing distillation methods that generalize across architectures is crucial for ensuring broad applicability. One promising direction is to distill representations that are invariant to specific network architectures. For example, future research could explore contrastive learning to enforce architecture-invariant feature alignment between CNNs and Transformers, ensuring that distilled datasets retain transferable representations. Another potential avenue is developing adversarial strategies to enforce feature robustness across different model families.
Towards Multimodal-Aware Dataset Distillation Future directions in dataset distillation should move beyond conventional image classification benchmarks and address more complex and multimodal domains, such as medical imaging, time-series forecasting, and multimodal sensor fusion. In these areas, it is essential to develop data representations tailored to specific modalities, such as spectral embeddings for audio signals or geometric descriptors for 3D point clouds [202]. Importantly, even when the final distilled data set remains unimodal (e.g. image only), recent studies [206] suggest that incorporating auxiliary multimodal knowledge, such as textual descriptions, segmentation masks, or audio cues, during the distillation process can significantly enhance the semantic richness and generalization ability of synthetic samples. Using multimodal supervision during training can significantly improve distillation quality across both vision and non-vision domains.
Robust and Fair Evaluation Frameworks To ensure fair benchmarking, future studies should decouple dataset informativeness from external enhancements like teacher distillation and data augmentation. Recent efforts such as DD-Ranking [207] introduce standardized evaluation metrics, but additional research is needed to refine benchmarking protocols and establish principled evaluation guidelines.
Strengthening Theoretical Foundations A rigorous theoretical framework is crucial for understanding dataset distillation’s limitations and guiding future advancements. Key challenges include quantifying information loss using entropy-based methods to establish lower bounds on minimal dataset requirements and applying PAC-learning [208] to derive generalization guarantees. Understanding the role of spectral representations in bridging architectural or distributional gaps remains critical. Recent work, such as UniDD’s spectral filtering framework [48], unifies existing methods into a spectral framework. Exploring optimal transport theory [209] for analyzing distributional shifts and integrating insights from information theory and statistical learning could also lead to more principled solutions. As dataset distillation continues to evolve, establishing a solid theoretical foundation will be essential for ensuring its scalability, generalization, and robustness across scenarios.
@@ Acknowledgment
The authors would like to express their gratitude to the contributors of the Awesome-Dataset-Distillation repository [210], which has served as a valuable resource in compiling recent advancements.